The Target Company is contemplating the replacement of their old printing machine with a new model costing $60,000. The old machine, which originally cost $40,00, has 6 years of expected life remaining and a current book value of $30,000 versus a current market value of $24,000. Targets corporate tax rate is 40%. If Target sells the old machine at market value, what is the initial outlay (after-tax) for the new printing machine?

22,000

30,000

33,000

-11,000
Answer #1

Book value of old machine = 30000

sale value = 24000

gain on sale = -6000

tax on gain = -6000 * 0.4 =-2400

After tax salvage value = 24000 - (-2400)

= 26400

Initial outlay machine = -60000 + 26400

= -33600